Solve z = (x – y)∅ (x² + y²)
1
Expert's answer
2020-05-05T18:33:52-0400

"z=(x-y)\phi(x^2+y^2)"

Let's denote "\partial z/\partial x=p, \,\, \partial z/\partial y=q".

Differentiate both sides with respect to "x" then to "y":

"p=(x-y) \phi'(x^2+y^2)2x+\phi(x^2+y^2)",

"q=(x-y) \phi'(x^2+y^2)2y-\phi(x^2+y^2)".

Hence, we have

"\frac{p-\phi(x^2+y^2)}{q+\phi(x^2+y^2)}=\frac{x}{y} \rightsquigarrow py-qx=\phi(x^2+y^2)(x+y)"